OpenStudy (doc.brown):

A skater rides off a drop of 35 degrees at 2.0m/s. When do they land on the decline?

OpenStudy (doc.brown):



OpenStudy (doc.brown):

The slope is tan(35) right?

OpenStudy (anonymous):

"when do they land on the decline" is that a 'time' question or a 'where on the incline do they land' question?

OpenStudy (anonymous):

@doc.brown

OpenStudy (anonymous):

 y-direction: s=s+vt+(1/2)at^2 y = 0 + 0 + (1/2)(-9.81)t^2 y=4.9t^2 [eq 1] x-direction s=s+vt+(1/2)at^2 x = 0 + 2(t) + (1/2) (0) t^2 x = 2t [eq 2] use angles given & combine eq 1 and eq 2 tan(35) = y/x tan(35) = ((-9.8/2) * t^2 ) / (2*t) t = 0.285 sec get values for x and y using eq 1 and 2 y=4.9t^2 [eq 1] y=4.9(0.285)^2 y=0.39m x = 2t [eq 2] x = 2(0.285) x = 0.57m find distance using pathagorean's therome x^2 + y^2 = d^2 0.57^2 + 0.39^2 = d^2 d = 0.69 d = 0.7 meter